Keith thinks of two numbers. When he adds the two numbers he gets eight. When he doubles the first number and takes away the treble the second number he gets 1.
Answer by josgarithmetic(39617)

x and y, the numbers.

x+y=8.

2x-3y=1.

I suggest elimination method to start.
-------------
2x+2y=16
-------------
2x-3y=1
-------------

Subtract as (2x+2y)-(2x-3y)=15
5y=15
y=3 and this means x=5.